A rectangle is three times as long as it is wide. If its perimeter is 48 cm, find its dimensions.
Answer by stanbon(75887)

A rectangle is three times as long as it is wide. If its perimeter is 48 cm, find its dimensions.
---------------
lenth = 3*width
-----
Perimeter = 2(length + width)
-------
48 = 2(3w + w)
24 = 4w
width = 6 cm
----
length = 3*6 = 18 cm
--------------
Cheers,
Stan H.
-------------------